OpenStudy (anonymous):

Perform the indicated operation(s) and simplify. Write answers with positive exponents and with integers or fractions (no decimals). (1/2x-2/3)(1/2x+2/3) (x^2+4x-8)divided by (2x)

OpenStudy (anonymous):

are these two separate expressions? first one is the difference of two squares (1/2x)^2 -(2/3)^2 = (1/4)x^(-2) - 4/9 second one is (x^2)/2x +4x/2x -8/2x = x/2 +4/2 -4/x